# LONGER: Scaling Up Long Sequence Modeling in Industrial Recommenders

### v1

@@ Abstract.
Modeling ultra-long user behavior sequences is critical for capturing both long- and short-term preferences in industrial recommender systems. Existing solutions typically rely on two-stage retrieval or indirect modeling paradigms, incuring upstream-downstream inconsistency and computational inefficiency. In this paper, we present LONGER, a Long-sequence Optimized traNsformer for GPU-Efficient Recommenders. LONGER incorporates (i) a global token mechanism for stabilizing attention over long contexts, (ii) a token merge module with lightweight InnerTransformers and hybrid attention strategy to reduce quadratic complexity, and (iii) a series of engineering optimizations, including training with mixed-precision and activation recomputation, KV cache serving, and the fully synchronous model training and serving framework for unified GPU-based dense and sparse parameter updates. LONGER consistently outperforms strong baselines in both offline metrics and online A/B testing in both advertising and e-commerce services at ByteDance, validating its consistent effectiveness and industrial-level scaling laws. Currently, LONGER has been fully deployed at more than 10 influential scenarios at ByteDance, serving billion users.
Long Sequence Modeling, Industrial Recommenders, Scaling Law
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Recommender systems
@@ 1. INTRODUCTION
In recommendation systems, ultra-long user historical behavior sequences comprehensively encapsulate both long-term and short-term user preferences(He et al., 2023; de Souza Pereira Moreira et al., 2021). While early sequential modeling architectures have been extensively studied and widely adopted in both academia and industry, their applications remain largely confined to short-sequence scenarios (sequence lengths of $10^{2}-10^{3}$). Fully modeling long sequences (length ¿ $10^{3}$) offers significant benefits for recommendation accuracy, diversity, and mitigates the information cocoons phenomena. However, due to the computational constraints, current industry $de\ facto$ practices for long-sequence modeling primarily adopt the following strategies:
@@
  • Pre-trained User Embeddings(Zhou et al., 2020b; Hou et al., 2022; Liu et al., 2021). In industry, it is a common practice to pre-train the entire ultra-long sequence in a source model and derive a condensed user embedding (UE), which can then be transferred to downstream recommendation models. Leveraging high-performance advances GPUs, this method support pre-training with sequence of up to $10^{3}$ length and multiple-layered transformers.
  • Memory-augmented Models. The multi-channel user interest memory network (MIMN) (Pi et al., 2019) offers a neural Turing machine and memory induction unit-based structure for user sequence memorizing, and large memory network (LMN)(Lu et al., 2025) presents a lightweight structure with product quantiztion-based decomposition. The memory augmented recommendation model (MARM) (Lv et al., 2024) proposes a memory-for-computation trade-off paradigm, which caches the intermediate results from computationally intensive modules.
While these strategies significantly improve computational efficiency, they inevitably sacrifice raw full-sequence information loss due to the upstream-downstream inconsistency or the indirect perception of the original ultra-long sequence, and thus these approaches essentially provide an intermediate stage in the evolution toward end-to-end long-sequence modeling.
Recently, the rapid advancement of large language models, exemplified by GPT (Radford et al., 2018), has established predictable scaling laws - empirical principles predicting performance improvements with increased model size, data volume, and computility. These scaling laws have recently guided innovations in recommendation systems. For example, HSTU (Zhai et al., 2024) consists of a stack of identical self-attention layers connected by residual connections for modeling long sequences, which shows better performance than vanilla Transformer architectures. Wukong(Zhang et al., 2024b) develops a stacked factorization machine and linear compression block based architecture for interaction, and validates the scaling laws in recommendation.
@@
To this end, we propose the Long-sequence Optimized traNsformer for GPU-Efficient Recommenders, i.e., LONGER. In the framework, we organize the sequence input as the global tokens and raw sequences, based on which an inner-transformer based token merge methodology is developed for effectively reducing computing budget. Besides, as there are generally much noise existed in users’ ultra long sequences, we utilize an efficient hybrid attention strategy for improving computational efficiency while maintaining model performance. Besides, to fully deploy LONGER at industrial level with billion-user scale, we present a series of engineering optimizations, including a fully synchronous training and serving framework with mixed-precision and activation recomputation, and a KV cache serving strategy. Overall, the contributions are mainly summarized as follows:
  • We present LONGER, a long-sequence optimized transformer structure for GPU-efficient recommenders. It presents an industrial GPU-efficient viewpoint by optimizing transformer structures and scales up user sequence modeling length to 5000 in an end-to-end manner.
  • LONGER sufficiently improves computational efficiency improvement through token merge and hybrid attention strategies, which reduce ~50% FLOPs and are validated to be almost lossless in performance. Besides, fully-optimized industrial training and serving framework is devised for further improved GPU computational efficiency and online deployments.
  • Thorough experiments are performance to validate the efficacy. Offline experiments on a billion-scale industrial dataset, and online A/B tests on two influential business scenarios at Douyin [^1] An influential short-video platform with billion user-scale: https://www.douyin.com/ are conducted to validate its performance. Currently, LONGER has been extensively developed in dozens of scenarios at ByteDance, affecting billion users.
@@ 2.1. Traditional Short-Sequence Modeling
To date, industrial recommendation systems predominantly adhere to the combined modeling paradigm of both sequence modeling and feature interaction(Zeng et al., 2024; Wang et al., 2017). Within the framework, sequence modeling has long played a pivotal role in depicting user preferences. Among the extensive researches, the influential milestone began from deep interest network (DIN) (Zhou et al., 2018). The subsequent approaches including DIEN(Zhou et al., 2019), CAN(Zhou et al., 2020a), etc. Besides, multi-domain (Chai et al., 2025; Chang et al., 2023b), multi-interest (Li et al., 2019; Chai et al., 2022), and sequence denosing methods (Shin et al., 2024; Chen et al., 2022) are extensively approached for different aspects in modeling user preferences. Noted that most of such sophisticated-designed structures are developed for short sequence modeling, while the long sequence modeling methods have later attracts increasing attention.
@@ 2.2. Long-Sequence Modeling
As has discussed in Introduction, the long-sequence modeling methods can be generally categorized into two-stage retrieval, pre-trained user embedding, and memory-augmented models. Overall, the retrieval-based and pre-trained methods belong to a two-stage strategy, and the memory-enhanced models generally require long-term training periods to accumulate hit rates within the memory slots. Recently, some efforts have been made to directly model long sequence(Zhai et al., 2024; Zivic et al., 2024; Zhang et al., 2024a). However, a GPU-efficient long sequence modeling has not been sufficiently explored in large-scaled industrial recommender systems.
@@ 3.1. Problem Statement
Let $\mathcal{U}$ and $\mathcal{I}$ denote the user and item sets respectively. Given a user $u\in\mathcal{U}$ with raw behavior sequence $S_{u}=[i^{(u)}_{1},...,i^{(u)}_{L}]$ where $i^{(u)}_{t}\in\mathcal{I}$, user basic features $u_{d}$ including user profiles, context features, and cross features, and a target item $v\in\mathcal{I}$, the recommendation task aims to predict the click or convert probability:
@@ 3.2. Overall Framework
To ensure scalability and deployment feasibility, we incorporate several engineering system-level optimizations. The framework provides fully synchronous training and serving with unified dense and sparse parameter storage on ultra-large-scaled GPU clusters. We further improve memory and compute efficiency with Mixed Precision Training and Recompute, enabling activation memory reduction and customized numerical precision. Finally, during inference, we deploy a KV Cache Serving strategy that caches user sequence representations and reuses them across candidate scoring, significantly reducing redundant computation.
@@ 3.3. Global Tokens
This architectural augmentation serves two primary purposes. First, global tokens function as centralized information anchors, enabling enhanced feature interactions between user history, contextual attributes, and candidate items. Second, they stabilize attention dynamics in long sequences, particularly under sparse attention configurations. As demonstrated in StreamLLM (Xiao et al., 2024), incorporating a small number of global tokens alleviates the “attention sink” effect, where deeper attention layers disproportionately focus on early tokens. Their presence acts as anchor points that maintain attention diversity and preserve long-range dependency modeling.
@@ 3.4. Token Merge
Let $L$ be sequence length and $d$ be embedding dimension. Processing long behavior sequences (typically $L\geq 2000$) with vanilla Transformers imposes prohibitive computational costs due to the quadratic attention complexity $O(L^{2}d)$, especially when $L\gg d$ (typically,  $L=2000$, $d=32$ in industrial recommenders). Conventional solutions such as sequence truncation lead to the loss of long-range dependencies. To address this, we propose a Token Merge strategy that groups adjacent tokens and compresses them into shorter sequences, achieving a trade-off between model efficiency and representational fidelity. This strategy reduces the sequence length by a factor of $K$, effectively performing spatial compression. The grouped token representations can be formed via simple concatenation or further enhanced by incorporating intra-group interactions through lightweight InnerTrans blocks. This design flexibly balances efficiency and expressiveness, preserving local semantics while enabling global modeling over a shorter sequence.
Given a standard-structured transformer encoder layer, the FLOPs and parameters can be expressed as:
@@
(4) #$\text{\#Params}_{\text{vanilla trans}}=12d^{2}+13d#$
@@ 3.5. LONGER Model Structure
In our model architecture, we use a hybrid attention mechanism that combines both cross attention and self-attention layers to efficiently process the input sequences.
@@ 3.5.1. Input Generation
3.5.2. Cross Causal Attention (First Layer)
In the first attention layer, we apply cross causal attention using the query matrix $\mathbf{O}$ generated in the previous step, and the input tokens $\mathbf{R}\in\mathbb{R}^{(m+L)\times d}$. The cross attention mechanism is computed as:
@@ 3.5.2. Cross Causal Attention (First Layer)
(10) $\mathbf{M}_{i,j}=\begin{cases}0,&\text{if }j\geq i,\ \text{where}\{i,j\}\in[1,d]\\ -\infty,&\text{otherwise}\end{cases}$
The causal mask design on one hand maintains temporal relevance between sequence items. On the other hand, it ensures the invisibility from the sequence to the candidate item, enabling the KV Cache Serving mechanism (see Section 3.6.3). After computing the attention, the result is passed through a feed-forward network (FFN) for further processing.
3.5.3. Self Causal Attention (Subsequent Layers)
After the cross causal attention layer, the subsequent layers consist of several self causal attention blocks. These layers focus on learning the internal relationships within the sampled tokens sequence, allowing the model to capture dependencies and patterns within the tokens of the behavior sequence itself. Each self causal attention layer is followed by an FFN, which helps in further processing the information learned by the attention mechanism. The self causal attention mechanism is computed using a similar formulation:
@@ 3.5.4. Stacking and Compression
The self causal attention layers are stacked $N$ times to iteratively refine the representations of the input sequence. After passing through these layers, the model produces a compressed output, which represents the final output of the attention mechanism. This output is then used for downstream prediction task.
@@
By using a combination of cross attention in the first layer and self attention in subsequent layers, our model is able to efficiently handle long sequences while leveraging both global context and internal dependencies.
@@ 3.6.2. Mixed Precision Training and Recompute
To alleviate GPU memory pressure during training, we adopt a recompute strategy alongside mixed precision training. For gradient computation, we use reverse-mode automatic differentiation, which is more efficient than forward-mode but requires storing all intermediate activations from the forward pass. These activations can become a major memory bottleneck. To address this, we support recompute declarations at the model definition level, allowing selected activations to be discarded during the forward pass and recomputed during the backward pass. This trades computation for memory savings. As native TensorFlow does not provide official support for recomputation, we implement it using the custom_gradient mechanism, enabling fine-grained control through code-level annotations.
@@ 4.1. Experimental Setting
For comparison, we evaluate our model against several strong baselines, categorized by their ability to model short- or long-range user behavior. Short-sequence methods include TWIN (Chang et al., 2023a) and DIN (Recent50) which rely on 50 interactions. Long-sequence methods, including SumPooling, DIN(Zhou et al., 2018), HSTU (Zhai et al., 2024), and Transformer(Chen et al., 2019), process extended behavior histories that often suffer from scalability and efficiency issues in industrial environments. All models are trained with the same preprocessing pipeline and hyperparameter tuning, and experiments are conducted on a 48×A100 GPU cluster.
@@ 4.2. Overall Performance
4.2.1. Comparison over exisisting methods.
We report model performance on the offline evaluation set using two standard metrics for binary classification in recommendation systems: AUC (Area Under the ROC Curve) and LogLoss. Table  1 summarizes the results across multiple baselines and our proposed model. According to the table, our model outperforms all baselines, achieving an AUC of 0.85290 and a LogLoss of 0.47103, which represents a relative improvement of 1.57% in AUC compared to the base model, and improves the AUC by 0.21% compared to the most competitive model, i.e., Transformer. It is noted that a 0.1% improvement is considered to be significant improvement that can affect the performance in online A/B test in the industrial case. Besides, the proposed model also show high efficiency compared to Transformer (see Section 4.2.2). This improvement demonstrates the effectiveness of our approach in capturing long-range behavior dependencies while maintaining computational efficiency.
@@
Overall, the ablation study confirms that both architectural enhancements (e.g. TokenMerge, InnerTrans) and query-related strategies (e.g., query number and selection method) play critical roles in balancing accuracy and efficiency. The findings validate that LONGER can achieve strong performance with reduced computational cost by carefully designing its key components and behavior modeling pipeline. Such a configuration makes our method highly suitable for large-scale industrial deployment, where low-latency inference and system throughput are essential.

### v2

@@ Abstract.
Modeling ultra-long user behavior sequences is critical for capturing both long- and short-term preferences in industrial recommender systems. Existing solutions typically rely on two-stage retrieval or indirect modeling paradigms, incurring upstream-downstream inconsistency and computational inefficiency. In this paper, we present LONGER, a Long-sequence Optimized traNsformer for GPU-Efficient Recommenders. LONGER incorporates (i) a global token mechanism for stabilizing attention over long contexts, (ii) a token merge module with lightweight InnerTransformers and hybrid attention strategy to reduce quadratic complexity, and (iii) a series of engineering optimizations, including training with mixed-precision and activation recomputation, KV cache serving, and the fully synchronous model training and serving framework for unified GPU-based dense and sparse parameter updates. LONGER consistently outperforms strong baselines in both offline metrics and online A/B testing in both advertising and e-commerce services at ByteDance, validating its consistent effectiveness and industrial-level scaling laws. Currently, LONGER has been validated and fully deployed across dozens of real-world influential scenarios at ByteDance, serving billions of users.
Ultra-Long Sequence Modeling, Industrial Recommenders, Scaling Law
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Nineteenth ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicbooktitle: Proceedings of the Nineteenth ACM Conference on Recommender Systems (RecSys ’25), September 22–26, 2025, Prague, Czech Republicdoi: 10.1145/3705328.3748065isbn: 979-8-4007-1364-4/2025/09ccs: Information systems Recommender systems
@@ 1. Introduction
In recommendation systems, ultra-long user historical behavior sequences comprehensively encapsulate both long-term and short-term user preferences(He et al., 2023; de Souza Pereira Moreira et al., 2021). While early sequential modeling architectures have been extensively studied and widely adopted in both academia and industry, their applications remain largely confined to short-sequence scenarios (sequence lengths of $10^{2}-10^{3}$). Fully modeling long sequences (length ¿ $10^{3}$) offers significant benefits for recommendation accuracy and diversity, and helps mitigate the information cocoon phenomenon. However, due to the computational constraints, current industry $de\ facto$ practices for long-sequence modeling primarily adopt the following strategies:
@@
  • Pre-trained User Embeddings(Zhou et al., 2020b; Hou et al., 2022; Liu et al., 2021). In industry, it is a common practice to pre-train the entire ultra-long sequence in a source model and derive a condensed user embedding (UE), which can then be transferred to downstream recommendation models. Leveraging high-performance advanced GPUs, this method supports pre-training with sequence of up to $10^{3}$ length and multiple-layered transformers.
  • Memory-augmented Models. The multi-channel user interest memory network (MIMN) (Pi et al., 2019) offers a neural Turing machine and memory induction unit-based structure for user sequence memorizing, and large memory network (LMN)(Lu et al., 2025) presents a lightweight structure with product quantization-based decomposition. The memory augmented recommendation model (MARM) (Lv et al., 2024) proposes a memory-for-computation trade-off paradigm, which caches the intermediate results from computationally intensive modules.
While these strategies significantly improve computational efficiency, they inevitably sacrifice raw full-sequence information due to the upstream-downstream inconsistency or the indirect perception of the original ultra-long sequence, and thus these approaches essentially provide an intermediate stage in the evolution toward end-to-end long-sequence modeling.
Recently, the rapid advancement of large language models, exemplified by GPT (Radford et al., 2018), has established scaling laws - empirical principles predicting performance improvements with increased model size, data volume, and computility. These scaling laws have recently guided innovations in recommendation systems. For example, HSTU (Zhai et al., 2024) consists of a stack of identical self-attention layers connected by residual connections for modeling long sequences, which shows better performance than vanilla Transformer architectures. Wukong(Zhang et al., 2024b) develops a stacked factorization machine and linear compression block based architecture for interaction, and validates the scaling laws in recommendation.
@@
To this end, we propose the Long-sequence Optimized traNsformer for GPU-Efficient Recommenders, i.e., LONGER. In the framework, we organize the sequence input as the global tokens and raw sequences, based on which an inner-transformer based token merge methodology is developed for effectively reducing computing budget. Besides, as there is generally a lot of noise present in users’ ultra long sequences, we utilize an efficient hybrid attention strategy for improving computational efficiency while maintaining model performance. Besides, to fully deploy LONGER at an industrial level with billion-user scale, we present a series of engineering optimizations, including a fully synchronous training and serving framework with mixed-precision and activation recomputation, and a KV cache serving strategy. Overall, the contributions are mainly summarized as follows:
  • We present LONGER, a long-sequence optimized transformer structure for GPU-efficient recommenders. It presents an industrial GPU-efficient viewpoint by optimizing transformer structures and scales up user sequence modeling length to 10,000 in an end-to-end manner in industry.
  • LONGER sufficiently improves computational efficiency through token merge and hybrid attention strategies, which reduce ~50% FLOPs and are validated to be almost lossless in performance. Besides, a fully-optimized industrial training and serving framework is devised for further improved GPU computational efficiency and online deployments.
  • Thorough experiments are conducted to validate the efficacy. Offline experiments on a billion-scale industrial dataset, and online A/B tests on two influential business scenarios at Douyin [^1] An influential short-video platform with billion user-scale: https://www.douyin.com/ are conducted to validate its performance. Currently, LONGER has been extensively developed in dozens of scenarios at ByteDance, affecting billions of users.
@@ 2.1. Traditional Short-Sequence Modeling
To date, industrial recommendation systems predominantly adhere to the combined modeling paradigm of both sequence modeling and feature interaction(Zeng et al., 2024; Wang et al., 2017). Within the framework, sequence modeling has long played a pivotal role in depicting user preferences. Among the extensive research, a pivotal milestone emerged with DIN (Zhou et al., 2018). The subsequent approaches including DIEN(Zhou et al., 2019), CAN(Zhou et al., 2020a), etc. Besides, multi-domain(Chai et al., 2025; Chang et al., 2023b), multi-interest(Li et al., 2019; Chai et al., 2022), and sequence denoising methods(Shin et al., 2024; Chen et al., 2022) are extensively approached for different aspects in modeling user preferences. Noted that most of such sophisticatedly designed structures are developed for short sequence modeling, while the long sequence modeling methods have later attracted increasing research attention.
@@ 2.2. Long-Sequence Modeling
As has been discussed in the Introduction, the long-sequence modeling methods can be generally categorized into two-stage retrieval, pre-trained user embedding, and memory-augmented models. Overall, the retrieval-based and pre-trained methods belong to a two-stage strategy, and the memory-enhanced models generally require long-term training periods to accumulate hit rates within the memory slots. Recently, some efforts have been made to directly model long sequence (Zhai et al., 2024; Zivic et al., 2024; Zhang et al., 2024a). However, a GPU-efficient long sequence modeling remains underexplored in large-scale industrial recommender systems.
@@ 3.1. Problem Statement
Let $\mathcal{U}$ and $\mathcal{I}$ denote the user and item sets, respectively. Given a user $u\in\mathcal{U}$ with raw behavior sequence $S_{u}=[i^{(u)}_{1},...,i^{(u)}_{L}]$ where $i^{(u)}_{t}\in\mathcal{I}$, user basic features $u_{d}$ including user profiles, context features, and cross features, and a target item $v\in\mathcal{I}$, the recommendation task aims to predict the click or convert probability:
@@ 3.2. Overall Framework
To ensure scalability and deployment feasibility, we incorporate several engineering system-level optimizations. The framework provides fully synchronous training and serving with unified dense and sparse parameter storage on ultra-large-scale GPU clusters. We further improve memory and compute efficiency with Mixed Precision Training and Recompute, enabling activation memory reduction and customized numerical precision. Finally, during inference, we deploy a KV Cache Serving strategy that caches user sequence representations and reuses them across candidate scoring, significantly reducing redundant computation.
@@ 3.3. Global Tokens
This architectural augmentation serves two primary purposes. First, global tokens function as centralized information anchors, enabling enhanced feature interactions between user history, contextual attributes, and candidate items. Second, they stabilize attention dynamics in long sequences, particularly under sparse attention configurations. As demonstrated in StreamLLM (Xiao et al., 2024), incorporating a small number of global tokens alleviates the “attention sink” effect, where deeper attention layers disproportionately focus on early tokens. These tokens act as anchor points that maintain attention diversity and preserve long-range dependency modeling.
@@ 3.4. Token Merge
Let $L$ be the sequence length and $d$ be the embedding dimension. Processing long behavior sequences (typically $L\geq 2000$) with vanilla Transformers imposes prohibitive computational costs due to the quadratic attention complexity $O(L^{2}d)$, especially when $L\gg d$ (typically,  $L=2000$, $d=32$ in industrial recommenders). Conventional solutions such as sequence truncation lead to the loss of long-range dependencies. To address this, we propose a Token Merge strategy that groups adjacent tokens and compresses them into shorter sequences, achieving a trade-off between model efficiency and representational fidelity. This strategy reduces the sequence length by a factor of $K$, effectively performing spatial compression. The grouped token representations can be formed via simple concatenation or further enhanced by incorporating intra-group interactions through lightweight InnerTrans blocks. This design provides a flexible trade-off between efficiency and expressiveness, preserving local semantics while enabling global modeling over a shorter sequence.
Given a standard-structured transformer encoder layer, the FLOPs and parameters can be expressed as (Narayanan et al., 2021):
@@
(4) $\text{Params}_{\text{vanilla trans}}=12d^{2}+13d$
@@ 3.5. LONGER Model Structure
In our model architecture, we use a hybrid attention mechanism that combines both cross-attention and self-attention layers to efficiently process the input sequences.
@@ 3.5.1. Input Generation
3.5.2. Cross-Causal Attention (First Layer)
In the first attention layer, we apply cross-causal attention using the query matrix $\mathbf{O}$ generated in the previous step, and the input tokens $\mathbf{R}\in\mathbb{R}^{(m+L)\times d}$. The cross-attention mechanism is computed as:
@@ 3.5.2. Cross-Causal Attention (First Layer)
(10) $\mathbf{M}_{i,j}=\begin{cases}0,&\text{if }j\geq i,\ \text{where}\{i,j\}\in[1,m+L]\\ -\infty,&\text{otherwise}\end{cases}$
The causal mask design, on one hand, maintains temporal relevance between sequence items. On the other hand, it ensures the invisibility from the sequence to the candidate item, enabling the KV Cache Serving mechanism (see Section 3.6.3). After computing the attention, the result is passed through a feed-forward network (FFN) for further processing.
3.5.3. Self-Causal Attention (Subsequent Layers)
After the cross-causal attention layer, the subsequent layers consist of several self-causal attention blocks. These layers focus on learning the internal relationships within the sampled tokens sequence, allowing the model to capture dependencies and patterns within the tokens of the behavior sequence itself. Each self-causal attention layer is followed by an FFN, which helps in further processing the information learned by the attention mechanism. The self-causal attention mechanism is computed using a similar formulation:
@@ 3.5.4. Stacking and Compression
The self-causal attention layers are stacked $N$ times to iteratively refine the representations of the input sequence. After passing through these layers, the model produces a compressed output, which represents the final output of the attention mechanism. This output is then used for the downstream prediction task.
@@
By using a combination of cross-attention in the first layer and self-attention in subsequent layers, our model is able to efficiently handle long sequences while leveraging both global context and internal dependencies.
@@ 3.6.2. Mixed Precision Training and Recompute
To alleviate GPU memory pressure during training, we adopt a recompute strategy alongside mixed precision training. For gradient computation, we use reverse-mode automatic differentiation, which is more efficient than forward-mode but requires storing all intermediate activations from the forward pass. These activations can become a major memory bottleneck. To address this, we support recomputing declarations at the model definition level, allowing selected activations to be discarded during the forward pass and recomputed during the backward pass. This trades computation for memory savings. As native TensorFlow does not provide official support for recomputation, we implement it using the custom_gradient mechanism, enabling fine-grained control through code-level annotations.
@@ 4.1. Experimental Setting
For comparison, we evaluate our model against several strong baselines, categorized by their ability to model short- or long-range user behavior. Short-sequence methods include TWIN (Chang et al., 2023a) and DIN (Recent50) which rely on 50 interactions. Long-sequence methods, including SumPooling, DIN(Zhou et al., 2018), HSTU (Zhai et al., 2024), and Transformer(Chen et al., 2019), process extended behavior histories that often suffer from scalability and efficiency issues in industrial environments. All models are trained with the same preprocessing pipeline and hyperparameter tuning, and experiments are conducted on a 48×A100s GPU cluster.
@@ 4.2. Overall Performance
4.2.1. Comparison of existing methods.
We report model performance on the offline evaluation set using two standard metrics for binary classification in recommendation systems: AUC (Area Under the ROC Curve) and LogLoss. Table  1 summarizes the results across multiple baselines and our proposed model. According to the table, our model outperforms all baselines, achieving an AUC of 0.85290 and a LogLoss of 0.47103, which represents a relative improvement of 1.57% in AUC compared to the base model, and improves the AUC by 0.21% compared to the most competitive model, i.e., Transformer. It is noted that a 0.1% improvement is considered to be a significant improvement that can affect the performance in online A/B test in the industrial case. Besides, the proposed model also demonstrates significantly higher efficiency compared to vanilla Transformer (see Section 4.2.2). This improvement demonstrates the effectiveness of our approach in capturing long-range behavior dependencies while maintaining computational efficiency.
@@
Overall, the ablation study confirms that both architectural enhancements (e.g., TokenMerge, InnerTrans) and query-related strategies (e.g., query number and selection method) play critical roles in balancing accuracy and efficiency. The findings validate that LONGER can achieve strong performance with reduced computational cost by carefully designing its key components and behavior modeling pipeline. Such a configuration makes our method highly suitable for large-scale industrial deployment, where low-latency inference and system throughput are essential.